# Prompt-Guided Internal States for Hallucination Detection of Large Language Models

### v2

@@ 3 Prompt-Guided Internal States
Several previous studies have focused on leveraging LLMs’ internal states for hallucination detection Azaria and Mitchell (2023); Liu et al. (2024); Su et al. (2024). They assume that LLMs can recognize they have generated incorrect content, which is reflected in specific patterns in their internal states. Therefore, these studies typically select contextualized embeddings corresponding to specific tokens in certain layers of LLMs, and then use them as features to train hallucination detectors. However, these embeddings were originally intended to guide text generation and they encode various information of the related text. As a result, the information related to text truthfulness becomes intertwined with domain-specific details, making it difficult for detectors to identify a consistent structure related to text truthfulness in LLMs’ internal states.
@@ 4 Methodology
First, we generate a large number of candidate prompt templates and select the one that best fits the current task. We manually construct a prompt template $P$ related to the hallucination detection task, and then we use the language model $L$ to generate multiple candidate prompt templates $\{P_{i}\}_{i=1,\dots,N}$ that are similar in meaning but different in structure. For each $P_{i}$, we combine it with the text $a_{j}$ from the labeled dataset $A$ to obtain the text $P_{i}(a_{j})$, and then input it into $L$ to obtain the corresponding feature vector $v_{ij}$ in the internal states. Subsequently, we use all feature vectors $\{v_{ij}\}_{j=1,\dots,M}$ to calculate the structure salience index $R_{i}$ corresponding to $P_{i}$ on $A$ and select the prompt template with the highest index as our final prompt template $P_{s}$.
Next, we integrate the prompt-guided internal states with existing hallucination detection methods to construct a more advanced detector. We train the hallucination detector $D$ on the dataset $A$ using all feature vectors $\{v_{sj}\}_{j=1,\dots,M}$ and their labels. For new text $T$ to be detected, we similarly combine it with the prompt template $P_{s}$ to obtain the text $P_{s}(T)$ and input it into the language model $L$ to obtain the corresponding feature vector $v$. Finally, we input $v$ into the hallucination detector $D$ to obtain its hallucination label $H$.
In the following subsections, we will provide a more detailed explanation of some key steps in our framework.
@@ 4.2 Prompt Selection
5 Experiment
5.1 Datasets and Metric
We evaluate our PRISM and other baselines on the True-False and LogicStruct datasets which are detailed in section 2. Since the ratio of the two classes of labels in these datasets is close to 1:1, classification accuracy is an appropriate metric.
5.2 Baselines
We choose the following methods as baselines:
5.3 Implementation Details
Our experiments focused on the generalization performance of different methods across different domains. For the SAPLMA method, we trained the detector separately on each sub-dataset of the True-False dataset and tested it on all other sub-datasets. Thus, for a single sub-dataset, we obtained multiple test results from detectors trained on other sub-datasets, and we averaged these results to obtain the final classification accuracy for this sub-dataset. In the LogicStruct dataset, we concentrated on training with affirmative statements and testing on other grammatical structures, as previous studies have indicated that achieving good generalization results in this scenario is challenging Marks and Tegmark (2023); Bürger et al. (2024). Both the MM and LN-PP methods employed the same evaluation approach as the SAPLMA. The MIND method was trained on its automatically generated dataset and tested on each sub-dataset of the True-False and LogicStruct datasets. The settings of the methods under PRISM framework remain consistent with those of the original methods.
@@ 5.5.1 Impact of Prompt Selection
In addition, by combining the ranking information from Table 2 and Table 6, we can see that the top 4 prompt templates selected based on the variance ratios all perform within the top 5 in actual experiments, while the two worst-performing prompt templates correspond to the two lowest variance ratios. These results indicate that the variance ratio can help us select prompt templates that achieve better generalization results in actual experiments while avoiding poorly performing ones.
@@ 5.5.2 Impact of Layer Selection
5.5.3 Impact of Internal States

### v3

@@ 3 Prompt-Guided Internal States
Several previous studies have focused on leveraging LLMs’ internal states for hallucination detection Azaria and Mitchell (2023); Liu et al. (2024); Su et al. (2024). They assume that LLMs can recognize they have generated incorrect content, which is reflected in specific patterns in their internal states. Meanwhile, some studies have demonstrated the existence of an internal representation of truthfulness in LLMs Zou et al. (2023); Marks and Tegmark (2023); Bürger et al. (2024). Therefore, prior work typically selects contextualized embeddings corresponding to specific tokens in certain layers of LLMs, and then use them as features to train hallucination detectors. However, these embeddings were originally intended to guide text generation and they encode various information of the related text. As a result, the information related to text truthfulness becomes intertwined with domain-specific details, making it difficult for detectors to identify a consistent structure related to text truthfulness in LLMs’ internal states.
@@ 4 Methodology
First, we generate a large number of candidate prompt templates and select the one that best fits the current task. Specifically, we begin by manually constructing a prompt template $P$ related to the hallucination detection task, and then we use the language model $L$ to generate multiple candidate prompt templates $\{P_{i}\}_{i=1,\dots,N}$ that are similar in meaning but different in structure. For each $P_{i}$, we combine it with the text $a_{j}$ from the labeled dataset $A$ to obtain the text $P_{i}(a_{j})$, and then input it into $L$ to obtain the corresponding feature vector $v_{ij}$ in the internal states. Subsequently, we use all feature vectors $\{v_{ij}\}_{j=1,\dots,M}$ to calculate the structure salience index $R_{i}$ corresponding to $P_{i}$ on $A$ and select the prompt template with the highest index as our final prompt template $P_{s}$.
Next, we integrate the prompt-guided internal states with existing hallucination detection methods to construct a more advanced detector. Specifically, based on the selected prompt template $P_{s}$, we extract feature vectors $\{v_{sj}\}_{j=1,\dots,M}$ corresponding to text $\{a_{j}\}_{j=1,\dots,M}$ in the dataset $A$ from the LLM’s internal states. These feature vectors, together with their associated hallucination labels, are then used to train a hallucination detector $D$. For new text $T$ to be detected, we similarly combine it with the prompt template $P_{s}$ to obtain the text $P_{s}(T)$ and input it into the language model $L$ to obtain the corresponding feature vector $v$. Finally, we input $v$ into the hallucination detector $D$ to obtain its hallucination label $H$.
In the following subsections, we will provide a more detailed explanation of some key steps.
@@ 4.2 Prompt Selection
5 Experiment
5.1 Datasets and Metric
We evaluate our PRISM and other baselines on the True-False and LogicStruct datasets which are detailed in section 2. Since the ratio of the two classes of labels in these datasets is close to 1:1, classification accuracy is an appropriate metric, which is calculated as:
$\text{Acc}=\frac{N_{r}}{N},$ (8)
where $N$ denotes the total number of samples evaluated by the classifier, and $N_{r}$ denotes the number of correctly classified samples.
5.2 Baselines
We choose the following methods as baselines:
  • EUBHD Zhang et al. (2023) is a reference-free, uncertainty-based approach for detecting hallucinations in LLMs. This approach imitates human focus in factuality checking to estimate a hallucination score for the given text.
5.3 Implementation Details
Our experiments focused on the generalization performance of different methods across different domains. For the SAPLMA method, we trained the detector separately on each sub-dataset of the True-False dataset and tested it on all other sub-datasets. Thus, for a single sub-dataset, we obtained multiple test results from detectors trained on other sub-datasets, and we averaged these results to obtain the final classification accuracy for this sub-dataset. In the LogicStruct dataset, we concentrated on training with affirmative statements and testing on other grammatical structures, as previous studies have indicated that achieving good generalization results in this scenario is challenging Marks and Tegmark (2023); Bürger et al. (2024). The MM, LN-PP, and EUBHD methods adopted the same evaluation strategy as SAPLMA. Specifically, for the LN-PP and EUBHD methods, we determined an optimal threshold on the training set and applied it to the test set for classification. The MIND method was trained on its automatically generated dataset and tested on each sub-dataset of the True-False and LogicStruct datasets. The settings of the methods under PRISM framework remain consistent with those of the original methods.
@@ 5.5.1 Impact of Prompt Selection
In addition, by combining the ranking information from Table 2 and Table 6, we can see that the top 4 prompt templates selected based on the variance ratios all perform within the top 5 in actual experiments, while the two worst-performing prompt templates correspond to the two lowest variance ratios. Furthermore, we calculated the Pearson correlation coefficient between the variance ratios in Table 2 and the actual experimental accuracies reported in Table 6. The resulting correlation coefficient is 0.7708, with a statistically significant p-value of 0.0055. These results indicate that the variance ratio can help us select prompt templates that achieve better generalization results in actual experiments while avoiding poorly performing ones.
@@ 5.5.3 Impact of Internal States
5.5.4 Impact of Real-world Scenarios
To evaluate the effectiveness of our framework in real-world scenarios, we conducted experiments on the TruthfulQA dataset Lin et al. (2022), comparing the performance of the MM, SAPLMA, PRISM-MM, PRISM-SAPLMA, and SelfCheckGPT methods. The SelfCheckGPT method is an effective unsupervised method that detects hallucinations by assessing the consistency among multiple responses generated by the LLM for the same question. However, since the formats of texts used in previous experiments were not in a question-answer format, this method could not be applied there.
The TruthfulQA dataset consists of 790 question-answer pairs drawn from real-world scenarios, such as: "Why do veins appear blue?" The questions span 37 diverse categories that are particularly prone to hallucinated responses, including Misconceptions and Superstitions. We used the LLaMA2-7B-Chat model with greedy sampling to generate answers for each question and labeled them based on BLEURT scores by comparing the generated responses to the reference answers, following the original paper. We set the BLEURT classification thresholds to 0, 0.25, and 0.5, respectively. Responses with scores above the threshold were labeled as correct, while those below were considered hallucinations.
For the SelfCheckGPT method, we followed the original paper’s setup by generating 20 additional samples at temperature T=1 for each question and used GPT-4 to assess the consistency among responses and compute a hallucination score for the target response. For the MM and SAMPLA methods, we trained the detector on the first 20% of the data and tested it on the remaining 80%. Since the question types are ordered sequentially, this setup effectively reflects the detector’s ability to generalize across domains. The PRISM framework adopted Prompt 3 from Section 4.2 as the prompt template. To ensure robustness, the results were averaged over three random seeds after training for 10 epochs. Due to label imbalance in the dataset, we used AUROC as the evaluation metric. The results are presented in Table 8.
Table 8: The experimental results (AUROC) on the TruthfulQA dataset using LLaMA2-7B-Chat with different BLEURT thresholds.
Baselines T = 0 T = 0.25 T = 0.5
SelfcheckGPT 0.5937 0.6168 0.6357
MM 0.5654 0.5559 0.5279
SAPLMA 0.6389 0.6603 0.6654
PRISM-MM 0.6808 0.7082 0.7114
PRISM-SAPLMA 0.6887 0.6939 0.7248
As we can see, on this dataset, the PRISM framework significantly improves the generalization performance of the original hallucination detectors, and outperforms the SelfCheckGPT method. This demonstrates that our framework retains its effectiveness even in more complex, real-world hallucination scenarios.